In a family with two children,find the mean of the number of children who will be girls

1
Expert's answer
2022-02-13T17:47:05-0500

Let "X="  the number of girls in a family with two children.


"P(X=0)=P(BB)=(\dfrac{1}{2})^2=\dfrac{1}{4}"

"P(X=1)=P(GB)+P(BG)=(\dfrac{1}{2})^2+(\dfrac{1}{2})^2=\dfrac{1}{2}"

"P(X=2)=P(GG)=(\dfrac{1}{2})^2=\dfrac{1}{4}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 x & 0 & 1 & 2 \\ \hline
 p(x) & 1/4 & 1/2 & 1/4 \\
 
\end{array}"

"mean=\mu=E(X)=\sum _ix_ip(x_i)"

"=0(\dfrac{1}{4})+1(\dfrac{1}{2})+2(\dfrac{1}{4})=1"